# Transfer Learning on Edge Connecting Probability Estimation under Graphon Model

### v1

@@ Appendix A Notation Table
Appendix C Additional Simulation Results
C.1 Additional results for Impact of Density Shift
@@
To evaluate the robustness of GTrans under density discrepancies between the source and target domains, we introduce a density shift parameter $\lambda\in[-0.5,0.5]$ to simulate structured perturbations, defined as $f_{s}(u,v)=f_{t}(u,v)+\xi$, where $\xi\sim U(0,\lambda)$ if $\lambda>0$ and $\xi\sim U(\lambda,0)$ if $\lambda<0$. We investigate the performance of GTrans-GW and GTrans-EGW across 10 different graphon functions (IDs: 1 to 10) with a fixed target sample size of 50 and a source sample size of 200. Each experiment is repeated 50 times to evaluate the adaptability of the models under varying levels of density shift. The results are summarized in Figure C1.
@@
Figure C1: MSE performance of five methods as the density shift parameter $\lambda$ varies from 0.5-0.5 to $0.5$, with error bars representing $\pm 0.1$ standard deviations.GTrans-GW (red circles with solid line), GTrans-EGW (pink circles with dashed line), NS (blue squares with solid line), USVT (yellow triangles with solid line), ICE (green diamonds with solid line), SAS (gray hollow squares with solid line). Top row: graphons 1-5; bottom row: graphons 5-10.
@@
Figure C1 shows the average MSE of different methods as the density shift parameter $\lambda$ varies from 0.5-0.5 to $0.5$. Both GTrans-GW (red solid line) and GTrans-EGW (pink dashed line) demonstrate substantial improvements over the baseline methods (NS, USVT, ICE, and SAS), which remain constant across all $\lambda$ values. Notably, as $\lambda$ shifts from negative to positive, GTrans methods effectively adjust their alignment, capturing structural changes and minimizing MSE. This result underscores the capability of GTrans-GW and GTrans-EGW to dynamically adapt to source-target discrepancies, achieving consistently lower MSE across all scenarios. Also, we notice that for dense graphons like Graphon 2 (density = 0.5) and Graphon 9 (density = 0.5), GTrans-GW and GTrans-EGW exhibit symmetric U-shaped MSE curves around zero, indicating balanced performance regardless of source density. In contrast, for sparse graphons like Graphon 6 (density = 0.13) and Graphon 10 (density = 0.19), the lowest MSE is achieved with a slightly denser source, highlighting the need for stronger initial signals when the target is sparse. Unique structural patterns in Graphon 5 (anti-diagonal) and Graphon 7 (oscillatory) cause non-monotonic fluctuations under density shift, reflecting sensitivity to structural misalignments.
C.2 Additional Simulation for Transfer Between Different Graphons
@@
Table C2: Comparison of different methods for various transfer scenarios. Best-performing methods are bolded.

### v2

@@ Appendix A Notation Table
Appendix C Proof of Theorem 4.2
Proof.
The proof is simple, which depends on the estimator error of $\mathbf{P}_{s}$ and $\mathbf{P}_{t}$. Recall that:
$\mathrm{GW}^{2}_{2}(\hat{\mathbf{P}}_{s}^{\rm init},\hat{\mathbf{P}}_{t}^{\rm init})$ $=\inf_{\pi\in\Gamma_{n}}\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}-\hat{P}_{T,jj^{\prime}}|^{2}\pi_{ij}\pi_{i^{\prime}j^{\prime}}$
$\mathrm{GW}^{2}_{2}(\mathbf{P}_{s},\mathbf{P}_{t})$ $=\inf_{\pi\in\Gamma_{n}}\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|P_{S,ii^{\prime}}-P_{T,jj^{\prime}}|^{2}\pi_{ij}\pi_{i^{\prime}j^{\prime}}$
where $\Gamma_{n}$ is the set of couplings, i.e.
$\Gamma_{n}=\left\{\pi:\sum_{i}\pi_{ij}=\frac{1}{n_{T}},\quad\sum_{j}\pi_{ij}=\frac{1}{n_{S}}\right\}\,.$
Now let $\hat{\pi}_{n}$ and $\pi_{n}$ denotes the minimizer/optimal coupling of $\mathrm{GW}^{2}_{2}(\hat{\mathbf{P}}_{s}^{\rm init},\hat{\mathbf{P}}_{t}^{\rm init})$ and $\mathrm{GW}^{2}_{2}(\mathbf{P}_{s},\mathbf{P}_{t})$ respectively. Then we have:
$\mathrm{GW}^{2}_{2}(\mathbf{P}_{s},\mathbf{P}_{t})$
$=\inf_{\pi\in\Gamma_{n}}\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\mathbf{P}_{S,ii^{\prime}}-\mathbf{P}_{T,jj^{\prime}}|^{2}\pi_{ij}\pi_{i^{\prime}j^{\prime}}$
$\leq\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\mathbf{P}_{S,ii^{\prime}}-\mathbf{P}_{T,jj^{\prime}}|^{2}\hat{\pi}_{n,ij}\hat{\pi}_{n,i^{\prime}j^{\prime}}$
$=\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\mathbf{P}_{S,ii^{\prime}}-\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}+\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}-\hat{\mathbf{P}}^{\rm init}_{T,jj^{\prime}}+\hat{\mathbf{P}}^{\rm init}_{T,jj^{\prime}}-\mathbf{P}_{T,jj^{\prime}}|^{2}\hat{\pi}_{n,ij}\hat{\pi}_{n,i^{\prime}j^{\prime}}$
$\leq 4\left[\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}-\mathbf{P}_{S,ii^{\prime}}|^{2}\hat{\pi}_{n,ij}\hat{\pi}_{n,i^{\prime}j^{\prime}}\right$.
$\qquad\left.+\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}-\hat{\mathbf{P}}^{\rm init}_{T,jj^{\prime}}|^{2}\hat{\pi}_{n,ij}\hat{\pi}_{n,i^{\prime}j^{\prime}}+\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\mathbf{P}_{T,jj^{\prime}}-\hat{\mathbf{P}}^{\rm init}_{T,jj^{\prime}}|^{2}\hat{\pi}_{n,ij}\hat{\pi}_{n,i^{\prime}j^{\prime}}\right]$
$\leq 4\left[\frac{1}{n_{S}^{2}}\sum_{i,i^{\prime}=1}^{n_{S}}|\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}-\mathbf{P}_{S,ii^{\prime}}|^{2}+\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}-\hat{\mathbf{P}}^{\rm init}_{T,jj^{\prime}}|^{2}\hat{\pi}_{n,ij}\hat{\pi}_{n,i^{\prime}j^{\prime}}+\frac{1}{n_{T}^{2}}\sum_{j,j^{\prime}=1}^{n_{T}}|\mathbf{P}_{T,jj^{\prime}}-\hat{\mathbf{P}}^{\rm init}_{T,jj^{\prime}}|^{2}\right]$
$\leq 4\left(\delta_{n_{S}}+\mathrm{GW}^{2}_{2}(\hat{\mathbf{P}}^{\rm init}_{S},\hat{\mathbf{P}}^{\rm init}_{T})+\delta_{n_{T}}\right)=4\left(\delta_{n}+\mathrm{GW}^{2}_{2}(\hat{\mathbf{P}}^{\rm init}_{S},\hat{\mathbf{P}}^{\rm init}_{T})\right)\,.$
Similarly, for the other side:
$\mathrm{GW}^{2}_{2}(\hat{\mathbf{P}}_{s}^{\rm init},\hat{\mathbf{P}}_{t}^{\rm init})$
$=\inf_{\pi\in\Gamma_{n}}\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}-\hat{\mathbf{P}}^{\rm init}_{T,jj^{\prime}}|^{2}\pi_{ij}\pi_{i^{\prime}j^{\prime}}$
$\leq\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}-\hat{\mathbf{P}}^{\rm init}_{T,jj^{\prime}}|^{2}\pi_{n,ij}\pi_{n,i^{\prime}j^{\prime}}$
$=\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}-\mathbf{P}_{S,ii^{\prime}}+\mathbf{P}_{S,ii^{\prime}}-\mathbf{P}_{T,jj^{\prime}}+\mathbf{P}_{T,jj^{\prime}}-\hat{\mathbf{P}}^{\rm init}_{T,jj^{\prime}}|^{2}\pi_{n,ij}\pi_{n,i^{\prime}j^{\prime}}$
$\leq 4\left[\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}-\mathbf{P}_{S,ii^{\prime}}|^{2}\pi_{n,ij}\pi_{n,i^{\prime}j^{\prime}}\right$.
$\qquad\left.+\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\mathbf{P}_{S,ii^{\prime}}-\mathbf{P}_{T,jj^{\prime}}|^{2}\pi_{n,ij}\pi_{n,i^{\prime}j^{\prime}}+\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\mathbf{P}_{T,jj^{\prime}}-\hat{\mathbf{P}}^{\rm init}_{T,jj^{\prime}}|^{2}\pi_{n,ij}\pi_{n,i^{\prime}j^{\prime}}\right]$
$\leq 4\left[\frac{1}{n_{S}^{2}}\sum_{i,i^{\prime}=1}^{n_{S}}|\hat{\mathbf{P}}^{\rm init}_{S,ii^{\prime}}-\mathbf{P}_{S,ii^{\prime}}|^{2}+\sum_{i,i^{\prime}=1}^{n_{S}}\sum_{j,j^{\prime}=1}^{n_{T}}|\mathbf{P}_{S,ii^{\prime}}-\mathbf{P}_{T,jj^{\prime}}|^{2}\pi_{n,ij}\pi_{n,i^{\prime}j^{\prime}}+\frac{1}{n_{T}^{2}}\sum_{j,j^{\prime}=1}^{n_{T}}|\mathbf{P}_{T,jj^{\prime}}-\hat{\mathbf{P}}^{\rm init}_{T,jj^{\prime}}|^{2}\right]$
$\leq 4\left(\delta_{n_{S}}+\mathrm{GW}^{2}_{2}(\mathbf{P}_{s},\mathbf{P}_{t})+\delta_{n_{T}}\right)=4\left(\delta_{n}+\mathrm{GW}^{2}_{2}(\mathbf{P}_{s},\mathbf{P}_{t})\right)\,.$
Appendix D Proof of Theorem 4.3
Proof.
We start with the definition of the GW distance:
$\mathrm{GW}_{2}^{2}(f,g)$ $=\int\int(f(x,x^{\prime})-g(y,y^{\prime}))^{2}\ d\pi^{*}(x,y)\ d\pi^{*}(x^{\prime},y^{\prime})$
$=\int f^{2}(x,x^{\prime})\ dx\ dx^{\prime}+\int g^{2}(y,y^{\prime})\ dy\ dy^{\prime}-2\int\int f(x,x^{\prime})g(y,y^{\prime})d\pi^{*}(x,y)\ d\pi^{*}(x^{\prime},y^{\prime})$
$=\int f^{2}(x,x^{\prime})\ dx\ dx^{\prime}+\int g^{2}(y,y^{\prime})\ dy\ dy^{\prime}-2\int\int\tilde{g}(y,y^{\prime})g(y,y^{\prime})\ dy\ dy^{\prime}$
Furthermore, we have:
$\int\tilde{g}^{2}(y,y^{\prime})\ dy\ dy^{\prime}$ $=\int\left(\int_{[0,1]^{2}}\pi^{*}(y,x)f(x,x^{\prime})\pi^{*}(x^{\prime},y)\ dx\ dx^{\prime}\right)^{2}\ dydy^{\prime}$
$\leq\int\int f^{2}(x,x^{\prime})\pi^{*}(y,x)\pi^{*}(x^{\prime},y)\ dxdx^{\prime}dydy^{\prime}$
$=\int f^{2}(x,x^{\prime})\ dx\ dx^{\prime}\,.$
Therefore, we have:
$\|\tilde{g}-g\|_{2}^{2}$ $=\int_{[0,1]^{2}}\ (\tilde{g}(y,y^{\prime})-g(y,y^{\prime}))^{2}\ dydy^{\prime}$
$=\int\tilde{g}^{2}(y,y^{\prime})\ dydy^{\prime}+\int g^{2}(y,y^{\prime})\ dydy^{\prime}-2\int\tilde{g}(y,y^{\prime})g(y,y^{\prime})\ dydy^{\prime}$
$\leq\int f^{2}(x,x^{\prime})\ dxdx^{\prime}+\int g^{2}(y,y^{\prime})\ dydy^{\prime}-2\int\int f(x,x^{\prime})g(y,y^{\prime})\ d\pi^{*}(x,y)d\pi^{*}(x^{\prime},y^{\prime})$
$=GW_{2}^{2}(f,g)\,.$
This completes the proof. ∎
Appendix E Additional Simulation Results
E.1 Additional results for Impact of Density Shift
@@
To evaluate the robustness of GTrans under density discrepancies between the source and target domains, we introduce a density shift parameter $\lambda\in[-0.5,0.5]$ to simulate structured perturbations, defined as $f_{s}(u,v)=f_{t}(u,v)+\xi$, where $\xi\sim U(0,\lambda)$ if $\lambda>0$ and $\xi\sim U(\lambda,0)$ if $\lambda<0$. We investigate the performance of GTrans-GW and GTrans-EGW across 10 different graphon functions (IDs: 1 to 10) with a fixed target sample size of 50 and a source sample size of 200. Each experiment is repeated 50 times to evaluate the adaptability of the models under varying levels of density shift. The results are summarized in Figure E1.
@@
Figure E1: MSE performance of five methods as the density shift parameter $\lambda$ varies from 0.5-0.5 to $0.5$, with error bars representing $\pm 0.1$ standard deviations.GTrans-GW (red circles with solid line), GTrans-EGW (pink circles with dashed line), NS (blue squares with solid line), USVT (yellow triangles with solid line), ICE (green diamonds with solid line), SAS (gray hollow squares with solid line). Top row: graphons 1-5; bottom row: graphons 5-10.
@@
Figure E1 shows the average MSE of different methods as the density shift parameter $\lambda$ varies from 0.5-0.5 to $0.5$. Both GTrans-GW (red solid line) and GTrans-EGW (pink dashed line) demonstrate substantial improvements over the baseline methods (NS, USVT, ICE, and SAS), which remain constant across all $\lambda$ values. Notably, as $\lambda$ shifts from negative to positive, GTrans methods effectively adjust their alignment, capturing structural changes and minimizing MSE. This result underscores the capability of GTrans-GW and GTrans-EGW to dynamically adapt to source-target discrepancies, achieving consistently lower MSE across all scenarios. Also, we notice that for dense graphons like Graphon 2 (density = 0.5) and Graphon 9 (density = 0.5), GTrans-GW and GTrans-EGW exhibit symmetric U-shaped MSE curves around zero, indicating balanced performance regardless of source density. In contrast, for sparse graphons like Graphon 6 (density = 0.13) and Graphon 10 (density = 0.19), the lowest MSE is achieved with a slightly denser source, highlighting the need for stronger initial signals when the target is sparse. Unique structural patterns in Graphon 5 (anti-diagonal) and Graphon 7 (oscillatory) cause non-monotonic fluctuations under density shift, reflecting sensitivity to structural misalignments.
E.2 Additional Simulation for Transfer Between Different Graphons
@@
Table E2: Comparison of different methods for various transfer scenarios. Best-performing methods are bolded.